OpenStudy (lacris):

I've been trying to simplify this polynomial stuff but I keep getting the wrong answer :( please help me, I will give you a shiny medal! (x-1)(x^2-4)+2(x-1)(x+2)=(x-1) p, then p =? I keep getting -12x+9, but that's not a possibility :( Please give me detailed instructions, thank you, I would like to learn this :)

OpenStudy (lacris):

\[(x-1)(x^2-4)+2(x-1)(x+2)=(x-1) p, then p =?\]

OpenStudy (mjdennis):

If I read this correctly, then there is a (x-1) factor in every term. So if you divide both sides by (x-1), they should cancel out and you just have to add the remaining terms. I see no sensible way to get (-12x+9), so that isn't the answer.

OpenStudy (lacris):

I just multiplied them together, I think that was the wrong idea, oops

OpenStudy (lacris):

\[\frac{ (x-1)(x^2-4)+2(x-1)(x+2)=(x-1) }{ x-1 }\]

OpenStudy (lacris):

so I cancel x-1 out? so i get x^2-4 + 2(x+2) = ??

OpenStudy (lacris):

\[ x^2-4 + 2(x+2) = which becomes 2x+5 \]

OpenStudy (lacris):

I mean 2x-4 (x^2-4)

OpenStudy (lacris):

so its x^2+2x = p??

OpenStudy (lacris):

x(x+2)

OpenStudy (agent0smith):

Looks right. Either should be fine.

OpenStudy (mjdennis):

That's what I got. Is it an option?

OpenStudy (lacris):

yes it is :DD

OpenStudy (lacris):

thank you so much for helping, it means a lot!